OpenStudy (anonymous):

Simplify 4x^2(4x)^-2

OpenStudy (anonymous):

Do you know what negative exponents do to a quantity?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

4x^2*(4x)^-2 =4x^2/(4x)^2 Negative exponents bring a quantity to the bottom; it makes it the denominator. =4x^2/16x^2 distribute the squared for the denominator =4/16 =1/4